OpenStudy (anonymous):

if f(x)=4x+1 and g(x)=x^2-3 find a.f(g(1)) b.g(f(4))

OpenStudy (anonymous):

a) f(g(1)) = f ( x^2 -3) = f (1^2 -3) = f(-2) = 4(-2) +1 = -7 b) g(f(4)) = g (4x+1) = g (4(4) +1) = g (17) = (17^2) - 3 = 286

OpenStudy (anonymous):

thank :D